1. A ball is drawn randomly from a jar that contains 6 red balls, 2 white balls, and 5 yellow balls. Find the probability of the given event.
a. The ball is not red.
b. The ball is not white.

Answer by math_tutor2020(3816)


Answers:
(a) 7/13
(b) 11/13

Reason

There are 6+2+5 = 13 balls total
Of which, 6 are red and 13-6 = 7 are not red
That's how we arrive at 7/13 to represent the probability of selecting a non-red ball (i.e. either white or yellow)

The probability of selecting a non-white ball follows a very similar idea.
There are 2 white balls and 13-2 = 11 non-white balls out of 13 total.